The answer to part B is not 7.00, and the Ka value for glycolic acid given by my textbook is 1.47x10^-4.Calculate the pH of the solution that results when 20.0 mL of 0.1700 M glycolic acid is (This problem requires values in your
Answer #1


pka of glcolic acid = -logKa

            = -log(1.47*10^-4)

                    = 3.83
a) after dilution

Concentration of glcolic acid = 20*0.17/45 = 0.0755 M

pH of glcolic acid = 1/2(pka-logC)

pka = 3.83 , C = 0.17 M

pH of glcolic acid = 1/2(3.83-log0.17)

                   = 2.3

b) no of mol of glcolic acid = M*V

            = 20*0.17 = 3.4 mmol

no of mol of NaOH added = M*V

                        = 25*0.136 = 3.4 mmol

Concentration of salt = 3.4/(20+25) = 0.0756 M

pH = 7+1/2(pka+logC)

pka = 3.83 , C = 0.0756 M

pH = 7+1/2(3.83+log0.0756)

   = 8.35

c) no of mol of glcolic acid = M*V

            = 20*0.17 = 3.4 mmol

no of mol of NaOH added = M*V

                        = 25*0.2 = 5 mmol

Concentration of excess NaOH = 5/(20+25) = 0.111 M

pOH = -log[OH-]

    = -log0.111

    = 0.955

pH = 14-0.955 = 13.045

d) no of mol of glcolic acid(acid) = M*V

            = 20*0.17 = 3.4 mmol

no of mol of sodium glcolate(salt) = M*V

                        = 25*0.2 = 5 mmol

pH = pka + log(salt/acid)

pka of glcolic acid = -logKa

            = -log(1.47*10^-4)

                    = 3.83
pH = 3.83+log(5/3.4)

   = 41